OpenStudy (anonymous):

Solve (z+ 7)2 = 3. Need help quick! Will MEDAL

OpenStudy (anonymous):

(z+7)^2=3

OpenStudy (anonymous):

\[(z+7)^2=3\]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

move the square to the other side

OpenStudy (anonymous):

Ok \[z+7=9\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

z+7=3^2?

OpenStudy (anonymous):

\[(z+7)^2=3\]

OpenStudy (anonymous):

That's what we started with

OpenStudy (anonymous):

\[z+7 = \sqrt{3}\]

OpenStudy (anonymous):

Ahh ok

OpenStudy (anonymous):

then subtract 7 from both sides?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So\[-7\]

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

\[z=\sqrt{3-7}\]

OpenStudy (anonymous):

\[-7\pm \sqrt{3}\]

OpenStudy (anonymous):

\[z= \sqrt{3} -7\]

OpenStudy (anonymous):

yeah that is correct

OpenStudy (anonymous):

Awesome thanks :D

OpenStudy (anonymous):

no problem